1- Forward price is the value of the forward contract.(true or false)

2- Calculate the present value of $100 in 5 years. Assume 6.1% interest rate with continuous compounding.

Round your answer to the nearest 2 decimal points. For example, if your answer is $12.345, then enter "12.35" in the answer box.
Answer #1

1)

False

The price of a forward contract is fixed, meaning that it does not change throughout the life cycle of the contract because the underlying will be purchased at a later date. We can consider the price of the forward contract “embedded” into the contract. The forward value is the opposite and fluctuates as the market conditions change.

2)

Present value = Future value / er*t

Present value = 100 / e5*0.061

Present value = 100 / e0.305

Present value = 100 / 1.356625

Present value = $73.71